OpenStudy (anonymous):

Lim sinx/x = 0 ? x--> -Infiniti

OpenStudy (anonymous):

Can you use l'Hospital's rule?

OpenStudy (anonymous):

Nope , our teacher wants us to use only rules which we have studied yet

OpenStudy (anonymous):

Use squeeze theorem.

OpenStudy (anonymous):

Ok let me try

OpenStudy (anonymous):

It's a finite quantity divided by infinity, what do you think?

OpenStudy (anonymous):

Oh i took the limit as x my mistake^^

OpenStudy (anonymous):

Or maybe 1 and -1 felt really brave one day and gave negative infinity a run for its money.

OpenStudy (anonymous):

Here is a hint: \(-1 \leq \sin(x)\leq 1\)

OpenStudy (anonymous):

So its 0

OpenStudy (anonymous):

I got that far ,then divide by x and substitute

OpenStudy (anonymous):

Either its 0 or im retarded

OpenStudy (anonymous):

Please dont tell me its the second

OpenStudy (anonymous):

Okay, since \(x\) will be negative, when you divide by \(x\) the inequalities flip.

OpenStudy (anonymous):

But yes, it does go to \(0\).

OpenStudy (anonymous):

Dont you mean divide1 by x

OpenStudy (anonymous):

You divide the entire inequality. All things are divided by \(x\).

OpenStudy (anonymous):

\[ \frac 1x \leq \frac{\sin(x)}x\leq -\frac 1x \]